# SCAN: Semantic Document Layout Analysis for Textual and Visual Retrieval-Augmented Generation

### v2

@@ 4.3 Visual RAG Evaluation Results
OHR-Bench. Table 8 presents the result of OHR-Bench in visual RAG. When applying our SCAN approach to divide original pages into semantic chunks and performing visual RAG on these chunks, we observed an overall improvement of 5.6 points compared to processing entire pages. We can see that SCAN approach is effective for every category. Especially, in the RO (reading order) task, our method achieves an impressive 18.4 points improvement. Recall that the RO task requires examining different paragraphs and articles to summarize answers. This demonstrates that dividing a page image into independent semantic chunks enables the system to retrieve only the relevant paragraphs, avoiding distractions from unrelated content on the same page.
BizMMRAG and Allganize. Table 9 presents the visual RAG results for Japanese document datasets. The findings demonstrate that on both the BizMMRAG and Allganize benchmarks, our SCAN methodology exhibits substantial accuracy improvements. Specifically, SCAN improves 10.4 points on BizMMRAG and 5.6 points on Allganize. This result also shows that our SCAN approach enables the VLM to achieve significantly enhanced performance in Japanese VQA.
@@ 4.4 Cost Comparison of VLM Text Conversion
DocLayout-YOLO YOLO11-X (56.9M) 9.9 11.3 22.4 59.6
Beehive RT-DETR-L (42.8M) 17.4 4.8 25.3 65.8
@@
We randomly sampled 10 images from OHR-Bench and used Qwen-2.5-VL-72B for text conversion. [^3] We used a vLLM vllm server running on four NVIDIA RTX PRO 6000 Blackwell GPUs for efficient batch processing. Table 10 reports the comparison between single-page processing and multiple semantic chunk processing. The results show that both the number of input tokens and output tokens increase when using multiple semantic chunks. This implies that applying SCAN with API-based models that charge by token usage may lead to higher monetary cost. However, despite the increase in token counts, the average processing time is reduced. We attribute this to the substantial decrease in the number of input tokens per request to the VLM, which lowers the cost of attention computation. Specifically, while the average number of input tokens per chunk is 1,320.4 in the single-page setting, it is reduced to 780.9 (=9683.1/12.4=9683.1/12.4) in the multiple-chunk setting. Thus, SCAN not only enables the extraction of richer textual information but also reduces computational overhead.
@@ 5 Conclusion
In this work, we study semantic document layout analysis for RAG. To the best of our knowledge, there is no negative societal impact in this research. All our training data consist of publicly available PDFs from the internet, which similarly present no ethical concerns. Our SCAN model aims to improve information extraction without introducing biases in the underlying content. We believe that improved document analysis can enhance the accessibility of information for users across different languages and document formats. While our system improves RAG capabilities, users should still be mindful of the general limitations of AI systems when relying on generated answers.
@@
1. Our SCAN model operates based on spatial image layout. In certain documents where contents that should logically form a single semantic chunk is physically separated in space and do not fit in a rectangular box, our current model cannot yet establish these connections. This limitation could potentially be addressed through an additional trainable reading order model coupled with a semantic box merging mechanism.
@@ Appendix B Annotation Instructions for Our Semantic Layout Dataset
Semantic units are always annotated with axis-aligned rectangles. Rectangles are required not to overlap and, taken together, should cover all content in the page. When a single rectangle would inevitably contain content from multiple topics, the region is split so that each rectangle corresponds to a single topic. For instance, in the top-right example in Figure 3, the bottom-right semantic box is topically related to the upper box, but they are annotated separately because a single rectangle would also include the bottom-left region, which belongs to a different topic.
@@ Appendix C Training Details of Our Object Detection Models
We fine-tuned two off-the-shelf object detection models: YOLO11-X and RT-DETR-X. We followed the default settings provided by the Ultralytics framework (version 8.3.28) [^4] https://github.com/ultralytics/ultralytics/releases/tag/v8.3.28 mostly, but we explicitly set or tuned some important hyper-parameters as shown in Table 12. For the YOLO11-X fine-tuning, we used 4 NVIDIA L40 GPUs (48GB), which took about 4 hours to finish 30 epochs. For the RT-DETR-X fine-tuning, we used 8 NVIDIA A100 GPUs (80GB), which took about 16 hours to finish 120 epochs.
@@ E.1 Environment
For text conversion with general VLMs, we used 8 NVIDIA L40S (48GB) GPUs with INTEL(R) XEON(R) GOLD 6548N CPU, and the details are as follows.

### v3

@@ 4.3 Visual RAG Evaluation Results
OHR-Bench. Table 8 presents the results of OHR-Bench in visual RAG. When applying our SCAN approach to divide original pages into semantic chunks and performing visual RAG on these chunks, we observed an overall improvement of 5.6 points compared to processing entire pages. We can see that the SCAN approach is effective for every category. Especially in the RO (reading order) task, our method achieves an impressive 18.4-point improvement. Recall that the RO task requires examining different paragraphs and articles to summarize answers. This demonstrates that dividing a page image into independent semantic chunks enables the system to retrieve only the relevant paragraphs, avoiding distractions from unrelated content on the same page.
BizMMRAG and Allganize. Table 9 presents the visual RAG results for Japanese document datasets. The findings demonstrate that on both the BizMMRAG and Allganize benchmarks, our SCAN methodology exhibits substantial accuracy improvements. Specifically, SCAN improves by 10.4 points on BizMMRAG and 5.6 points on Allganize. This result also shows that our SCAN approach enables the VLM to achieve significantly enhanced performance in Japanese VQA.
@@ 4.4 Cost Comparison of VLM Text Conversion
DocLayout-YOLO YOLO11-X (57M) 9.9 11.3 22.4 59.6
Beehive RT-DETR-L (43M) 17.4 4.8 25.3 65.8
@@
We randomly sampled 10 images from OHR-Bench and used Qwen2.5-VL-72B for text conversion. [^3] We used a vLLM Kwon et al. (2023) server running on four NVIDIA RTX PRO 6000 Blackwell GPUs for efficient batch processing. Table 10 reports the comparison between single-page processing and multiple semantic chunk processing. The results show that both the number of input tokens and output tokens increase when using multiple semantic chunks. This implies that applying SCAN with API-based models that charge by token usage may lead to higher cost. However, despite the increase in token counts, the average processing time is reduced. We attribute this to the substantial decrease in the number of input tokens per request to the VLM, which lowers the cost of attention computation. Specifically, while the average number of input tokens per chunk is 1,320.4 in the single-page setting, it is reduced to 780.9 (=9683.1/12.4=9683.1/12.4) in the multiple-chunk setting. Thus, SCAN not only enables the extraction of richer textual information but also reduces computational overhead.
@@ 5 Conclusion
In this work, we study semantic document layout analysis for RAG. To the best of our knowledge, there is no negative societal impact in this research. All our training data consist of publicly available PDFs from the internet, which likewise present no ethical concerns. Our SCAN model aims to improve information extraction without introducing biases in the underlying content. We believe that improved document analysis can enhance the accessibility of information for users across different languages and document formats. While our system improves RAG capabilities, users should still be mindful of the general limitations of AI systems when relying on generated answers.
@@
1. Our SCAN model operates based on spatial image layout. In certain documents where content that should logically form a single semantic chunk is physically separated in space and does not fit in a rectangular box, our current model cannot yet establish these connections. This limitation could potentially be addressed through an additional trainable reading order model coupled with a semantic box merging mechanism.
@@ Appendix B Annotation Instructions for Our Semantic Layout Dataset
Semantic units are always annotated with axis-aligned rectangles. Rectangles are required not to overlap and, taken together, should cover all content on the page. When a single rectangle would inevitably contain content from multiple topics, the region is split so that each rectangle corresponds to a single topic. For instance, in the top-right example in Figure 3, the bottom-right semantic box is topically related to the upper box, but they are annotated separately because a single rectangle would also include the bottom-left region, which belongs to a different topic.
@@ Appendix C Training Details of Our Object Detection Models
We fine-tuned two off-the-shelf object detection models: YOLO11-X and RT-DETR-X. We mostly followed the default settings provided by the Ultralytics framework (version 8.3.28) [^4] https://github.com/ultralytics/ultralytics/releases/tag/v8.3.28, but we explicitly set or tuned some important hyper-parameters as shown in Table 12. For the YOLO11-X fine-tuning, we used 4 NVIDIA L40 GPUs (48GB), which took about 4 hours to finish 30 epochs. For the RT-DETR-X fine-tuning, we used 8 NVIDIA A100 GPUs (80GB), which took about 16 hours to finish 120 epochs.
@@ E.1 Environment
For text conversion with Qwen2.5-VL-72B, we used 8 NVIDIA L40S (48GB) GPUs and an INTEL(R) XEON(R) GOLD 6548N CPU. The details are as follows.